what is the number? it is an even, two digit number, the difference between the digit is 3, the one's digit is twice the ten's digit

Answer by lwsshak3(11628)

what is the number? it is an even, two digit number, the difference between the digit is 3, the one's digit is twice the ten's digit.
***
let u=units digit
let t=tens digit
u-t=3
t=u-3
u=2t
u=2(u-3)
u=2u-6
u=6
t=u-3=3
number=36